# matrix,arrows,decorations.pathmorphing

### v2

@@ 2.9. A rigidity criterion
Let us then denote by $\operatorname{Out}^{\sharp}(\mathrm{P}{\check{\Gamma}}(S_{0,n}))$ the centralizer of the image of $\Sigma_{n}$ in $\operatorname{Out}^{\mathbb{I}}(\mathrm{P}{\check{\Gamma}}(S_{0,n}))$. Since $\Sigma_{n}$ acts transitively on the set of conjugacy classes of procyclic subgroups of $\mathrm{P}{\check{\Gamma}}(S_{0,n})$ generated by Dehn twists about simple closed curves bounding $2$-punctured discs in $S_{0,n}$, it follows that $\operatorname{Out}^{\sharp}(\mathrm{P}{\check{\Gamma}}(S_{0,n}))$ preserves such conjugacy classes. Hence, for $n\leq 5$, our definition of this group agrees with the one given in Section 0.1 of [10].
@@
From the isomorphism $\mathrm{P}\Gamma(S_{1,2})\cong{\Gamma}(S_{0,5})_{Q}$ (cf. Section 2.4), it follows that there is a natural isomorphism $\operatorname{Inn}(\mathrm{P}\Gamma(S_{1,2}))\left/\operatorname{Inn}(\mathrm{P}{\check{\Gamma}}(S_{0,5}))\right.\cong\Sigma_{4}$. Let us then define $\operatorname{Out}^{\sharp}(\mathrm{P}{\check{\Gamma}}(S_{1,2}))$ to be the centralizer of the image of $\Sigma_{4}$ in $\operatorname{Aut}^{\mathbb{I}}(\mathrm{P}{\check{\Gamma}}(S_{1,2}))\left/\operatorname{Inn}(\mathrm{P}{\check{\Gamma}}(S_{0,5}))\right.$.
@@
For $S$ one of the surfaces considered above, let also $\operatorname{Aut}^{\sharp}(\mathrm{P}{\check{\Gamma}}(S))$ be the inverse image of $\operatorname{Out}^{\sharp}(\mathrm{P}{\check{\Gamma}}(S))$ in $\operatorname{Aut}(\mathrm{P}{\check{\Gamma}}(S))$ (note that $\operatorname{Aut}^{\sharp}(\mathrm{P}{\check{\Gamma}}(S))$ is contained in $\operatorname{Aut}^{\mathbb{I}}(\mathrm{P}{\check{\Gamma}}(S))$). We then have:
@@
$\begin{array}[]{l}\operatorname{Aut}^{\mathbb{I}}(\mathrm{P}{\check{\Gamma}}(S_{1,2}))\left/\operatorname{Inn}(\mathrm{P}{\check{\Gamma}}(S_{0,5}))\right.\cong\Sigma_{4}\times\operatorname{Out}^{\sharp}(\mathrm{P}{\check{\Gamma}}(S_{1,2})),\\ \\ \operatorname{Out}^{\sharp}(\mathrm{P}{\check{\Gamma}}(S_{1,2}))\cong\operatorname{Out}^{\mathbb{I}}(\mathrm{P}{\check{\Gamma}}(S_{1,2})),\\ \\ \operatorname{Out}^{\mathbb{I}}(\mathrm{P}{\check{\Gamma}}(S_{0,5}))\cong\Sigma_{5}\times\operatorname{Out}^{\sharp}(\mathrm{P}{\check{\Gamma}}(S_{0,5})).\end{array}$
@@
$\operatorname{Aut}^{\sharp}(\mathrm{P}{\check{\Gamma}}(S_{1,2}))\hookrightarrow\operatorname{Aut}^{\sharp}(\mathrm{P}{\check{\Gamma}}(S_{0,5})).$
@@
There are short exact sequences:
@@
and (cf. Corollary C in [11])
@@
Since $\Sigma_{4}$ and $\Sigma_{5}$ are complete groups, the above short exact sequences split and there are natural isomorphisms as stated in the lemma (cf. Theorem 7.15 in [23]). The last statement of the lemma follows from the fact that the centralizer of $\Sigma_{4}$ in $\Sigma_{5}$ is trivial. ∎
Remark 2.18.
By Corollary C in [20] and Main Theorem in [10], there is actually a series of natural isomorphisms $\operatorname{Out}^{\sharp}(\mathrm{P}{\check{\Gamma}}(S_{1,2}))\cong\operatorname{Out}^{\sharp}(\mathrm{P}{\check{\Gamma}}(S_{0,5}))\cong\widehat{\operatorname{GT}}$, where $\widehat{\operatorname{GT}}$ is the profinite Grothendieck-Teichmüller group, so that the monomorphism of Lemma 2.17 is actually an isomorphism. But we will not need this fact.
@@
If the given $\{v_{0},v_{1}\}\in{\check{C}}_{P}(S_{1,2})_{1}$ is an edge contained in a ${\check{\Gamma}}(S_{1,2})$-orbit such that the pair of vertices $\{q(v_{0}),q(v_{1})\}$ is an edge of ${\check{C}}_{P}(S_{0,5})$, then, since $f\in\operatorname{Aut}^{\mathbb{I}}(\mathrm{P}{\check{\Gamma}}(S_{1,2}))$, by Lemma 2.16, preserves ${\check{\Gamma}}(S_{1,2})$-orbits, we have that $\{\tilde{f}(q(v_{0})),\tilde{f}(q(v_{1}))\}$ is also an edge of ${\check{C}}_{P}(S_{0,5})$. In this case, we just let $w_{0}:=q(v_{0})$ and $w_{1}:=q(v_{1})$ and we are done. In particular, as it is easy to check, this happens if the common profinite simple closed curve in the intersection $v_{0}\cap v_{1}$ has the topological type of a separating curve on $S_{1,2}$.
@@
Let us then consider the case when the common profinite simple closed curve ${\gamma}$ in $v_{0}\cap v_{1}$ has the topological type of a nonseparating curve on $S_{1,2}$. As usual, it is not restrictive to assume that ${\gamma}\in C(S_{1,2})_{0}\subset{\check{C}}(S_{1,2})_{0}$ and after, possibly, composing the given $f\in\operatorname{Aut}^{\mathbb{I}}(\mathrm{P}{\check{\Gamma}}(S_{1,2}))$ with an inner automorphism of $\mathrm{P}\Gamma(S_{1,2})$, we can also assume that $f\in\operatorname{Aut}^{\sharp}(\mathrm{P}{\check{\Gamma}}(S_{1,2}))$. From Lemma 2.17, it then follows that that the conjugacy class of ${\hat{\mathrm{I}}}_{\gamma}\cap\mathrm{P}{\check{\Gamma}}(S_{0,5})$ in $\mathrm{P}{\check{\Gamma}}(S_{0,5})$ is preserved by $\tilde{f}$. Therefore, after, possibly, composing by an inner automorphism of $\mathrm{P}{\check{\Gamma}}(S_{0,5})$, we can at last assume that $f\in\operatorname{Aut}^{\sharp}(\mathrm{P}{\check{\Gamma}}(S_{1,2}))_{{\hat{\mathrm{I}}}_{\gamma}}$.
@@
With the above assumptions, in particular, $f$ preserves the link $\operatorname{Lk}({\gamma})$ of ${\gamma}$ in ${\check{C}}(S_{1,2})$ and then the vertex set of the profinite subgraph $L_{\gamma}$ of ${\check{C}}_{P}(S_{1,2})$. Let us recall (cf. Lemma 2.8) that there are natural $\mathrm{P}{\check{\Gamma}}(S_{1,2})_{\gamma}$-equivariant continuous isomorphisms $\operatorname{Lk}({\gamma})\cong{\check{C}}(S_{1,2}\smallsetminus{\gamma})$ and $L_{\gamma}\cong{\check{C}}_{P}(S_{1,2}\smallsetminus{\gamma})$.
By hypothesis, we have that both $\{v_{0},v_{1}\}$ and $\{f(v_{0}),f(v_{1})\}\in(L_{\gamma})_{1}$. We claim that $f$ preserves all the edge set of $L_{\gamma}$ (and so induces an automorphism of this profinite graph). By Lemma 2.10, $f$ acts on the vertex set $(L_{\gamma})_{0}\cong{\check{C}}_{P}(S_{1,2}\smallsetminus{\gamma})_{0}$ through its image via the natural homomorphism $\operatorname{Aut}^{\mathbb{I}}(\mathrm{P}{\check{\Gamma}}(S_{1,2}))_{{\hat{\mathrm{I}}}_{\gamma}}\to\operatorname{Aut}^{\mathbb{I}}(\mathrm{P}{\check{\Gamma}}(S_{1,2}\smallsetminus{\gamma}))$. This is induced by the restriction to the stabilizer $\mathrm{P}{\check{\Gamma}}(S_{1,2})_{\vec{\gamma}}$ followed by the projection to its quotient $\mathrm{P}{\check{\Gamma}}(S_{1,2}\smallsetminus{\gamma})$.
By Lemma 2.17, there is a natural homomorphism $\operatorname{Aut}^{\sharp}(\mathrm{P}{\check{\Gamma}}(S_{1,2}))_{{\hat{\mathrm{I}}}_{\gamma}}\hookrightarrow\operatorname{Aut}^{\sharp}(\mathrm{P}{\check{\Gamma}}(S_{0,5}))$, induced by restriction. Since $\mathrm{P}{\check{\Gamma}}(S_{1,2})_{\vec{\gamma}}\cap\mathrm{P}{\check{\Gamma}}(S_{0,5})=\mathrm{P}{\check{\Gamma}}(S_{0,5})_{q({\gamma})}$ and the natural projection $\mathrm{P}{\check{\Gamma}}(S_{0,5})_{q({\gamma})}\to\mathrm{P}{\check{\Gamma}}(S_{0,5}\smallsetminus q({\gamma}))$, after identifying $\mathrm{P}{\check{\Gamma}}(S_{0,5}\smallsetminus q({\gamma}))$ with $\mathrm{P}{\check{\Gamma}}(S_{0,4})$, is just the restriction of a forgetful homomorphism $\mathrm{P}{\check{\Gamma}}(S_{0,5})\to\mathrm{P}{\check{\Gamma}}(S_{0,4})$ to $\mathrm{P}{\check{\Gamma}}(S_{0,5})_{q({\gamma})}$, we see that the restriction of the homomorphism $\operatorname{Aut}^{\mathbb{I}}(\mathrm{P}{\check{\Gamma}}(S_{1,2}))_{{\hat{\mathrm{I}}}_{\gamma}}\to\operatorname{Aut}^{\mathbb{I}}(\mathrm{P}{\check{\Gamma}}(S_{1,2}\smallsetminus{\gamma}))$ to the subgroup $\operatorname{Aut}^{\sharp}(\mathrm{P}{\check{\Gamma}}(S_{1,2}))_{{\hat{\mathrm{I}}}_{\gamma}}$ is equivalent to the restriction of the homomorphism $\operatorname{Aut}^{\sharp}(\mathrm{P}{\check{\Gamma}}(S_{0,5}))\to\operatorname{Aut}^{\sharp}(\mathrm{P}{\check{\Gamma}}(S_{0,4}))$, defined in Section 2.2 in [10], associated to the forgetful homomorphism considered above, to the image of $\operatorname{Aut}^{\sharp}(\mathrm{P}{\check{\Gamma}}(S_{1,2}))_{{\hat{\mathrm{I}}}_{\gamma}}$ in $\operatorname{Aut}^{\sharp}(\mathrm{P}{\check{\Gamma}}(S_{0,5}))$.
In particular, the image of $f$ in $\operatorname{Out}^{\mathbb{I}}(\mathrm{P}{\check{\Gamma}}(S_{1,2}\smallsetminus{\gamma}))$ commutes with the image of $\Sigma_{4}$ via the outer representation associated to the short exact sequence:
@@
$1\to\mathrm{P}{\check{\Gamma}}(S_{1,2}\smallsetminus{\gamma})\to{\check{\Gamma}}(S_{1,2}\smallsetminus{\gamma})\to\Sigma_{4}\to 1$
@@
It is then clear that all hypotheses of Lemma 2.11 are satisfied and $f$ extends to an automorphism of ${\check{\Gamma}}(S_{1,2}\smallsetminus{\gamma})$. As in Section 2.10, we then conclude that $f$ induces an automorphism of the pants complex ${\check{C}}_{P}(S_{1,2}\smallsetminus{\gamma})$, as claimed above.
@@
It is now easy to check that, for some edge $\{v_{0}^{\prime},v_{1}^{\prime}\}\in(L_{\gamma})_{1}\cong{\check{C}}_{P}(S_{1,2}\smallsetminus{\gamma})_{1}$, we have that $\{q(v_{0}^{\prime}),q(v_{1}^{\prime})\}\in{\check{C}}_{P}(S_{0,5})_{1}$ and conclude, as we did at the beginning of the proof, letting $w_{0}:=q(v_{0}^{\prime})$ and $w_{1}:=q(v_{1}^{\prime})$. ∎
@@
From Lemma 9.13 in [6], it now follows that the normalizer of $\mathrm{P}{\check{\Gamma}}(S_{1,2})$ in ${\check{\Gamma}}^{\pm}(S_{0,5})$ coincides with the closure in this group of the normalizer of $\mathrm{P}\Gamma(S_{1,2})$ in ${\Gamma}^{\pm}(S_{0,5})$, which, as it easily follows from (ii) of Theorem in [17], is just $\mathrm{P}\Gamma^{\pm}(S_{1,2})$. Since $\operatorname{Inn}({\check{\Gamma}}^{\pm}(S_{1,2}))=\operatorname{Inn}(\mathrm{P}{\check{\Gamma}}^{\pm}(S_{1,2}))$, this implies Theorem 2.15 for $S=S_{1,2}$.

### v3

@@ 2.9. A rigidity criterion
Let us denote by $\operatorname{Out}^{\mathbb{I}}_{\Sigma_{n}}(\mathrm{P}{\check{\Gamma}}(S_{0,n}))$ the centralizer of the image of $\Sigma_{n}$ in $\operatorname{Out}^{\mathbb{I}}(\mathrm{P}{\check{\Gamma}}(S_{0,n}))$. Let us recall that $\operatorname{Out}^{\sharp}(\mathrm{P}{\check{\Gamma}}(S_{0,n}))$ was defined in [16, Section 0.1], for $n\geq 4$, to be the subgroup of $\operatorname{Out}(\mathrm{P}{\check{\Gamma}}(S_{0,n}))$ consisting of those elements which commute with the image of $\Sigma_{n}$ and preserve each conjugacy class of the procyclic subgroups of $\mathrm{P}{\check{\Gamma}}(S_{0,n})$ generated by a Dehn twist about a simple closed curve bounding a $2$-punctured disc in $S_{0,n}$. We have:
Lemma 2.20.
For $n=4$ and $5$, there holds $\operatorname{Out}^{\sharp}(\mathrm{P}{\check{\Gamma}}(S_{0,n}))=\operatorname{Out}^{\mathbb{I}}_{\Sigma_{n}}(\mathrm{P}{\check{\Gamma}}(S_{0,n}))$.
Proof.
For $n=4,5$, all essential simple closed curves on $S_{0,n}$ bound a $2$-punctured disc. Therefore, there is at least an inclusion $\operatorname{Out}^{\sharp}(\mathrm{P}{\check{\Gamma}}(S_{0,n}))\subseteq\operatorname{Out}^{\mathbb{I}}_{\Sigma_{n}}(\mathrm{P}{\check{\Gamma}}(S_{0,n}))$. To prove that the reverse inclusion holds, we have to show that the elements of $\operatorname{Out}^{\mathbb{I}}_{\Sigma_{n}}(\mathrm{P}{\check{\Gamma}}(S_{0,n}))$ preserve every conjugacy class of procyclic subgroup of $\mathrm{P}{\check{\Gamma}}(S_{0,n})$ generated by a Dehn twist. By definition, the group $\operatorname{Out}^{\mathbb{I}}(\mathrm{P}{\check{\Gamma}}(S_{0,n}))$ acts on the set of such conjugacy classes with its subgroup $\Sigma_{n}$ acting transitively on it. This easily implies that those elements which commute with the action of $\Sigma_{n}$ act trivially on this set. ∎
@@
From the isomorphism $\mathrm{P}\Gamma(S_{1,2})\cong{\Gamma}(S_{0,5})_{Q}$ (cf. Section 2.4) and the isomorphisms (9), it follows that there is a series of isomorphisms $\operatorname{Inn}(\mathrm{P}\Gamma(S_{1,2}))\left/\operatorname{Inn}(\mathrm{P}{\check{\Gamma}}(S_{0,5}))\right.\cong(\Sigma_{5})_{Q}\cong\Sigma_{4}$. Let us then define:
  • $\widetilde{\operatorname{Out}}^{\mathbb{I}}(\mathrm{P}{\check{\Gamma}}(S_{1,2})):=\operatorname{Aut}^{\mathbb{I}}(\mathrm{P}{\check{\Gamma}}(S_{1,2}))\left/\operatorname{Inn}(\mathrm{P}{\check{\Gamma}}(S_{0,5}))\right$.;
  • $\widetilde{\operatorname{Out}}^{\mathbb{I}}_{{\Sigma}_{4}}(\mathrm{P}{\check{\Gamma}}(S_{1,2}))$ to be the centralizer of the image of $\Sigma_{4}$ in $\widetilde{\operatorname{Out}}^{\mathbb{I}}(\mathrm{P}{\check{\Gamma}}(S_{1,2}))$;
  • $\operatorname{Aut}^{\mathbb{I}}_{{\Sigma}_{5}}(\mathrm{P}{\check{\Gamma}}(S_{0,5}))$ (resp. $\operatorname{Aut}^{\mathbb{I}}_{{\Sigma}_{4}}(\mathrm{P}{\check{\Gamma}}(S_{1,2}))$) to be the inverse image of $\operatorname{Out}^{\mathbb{I}}_{{\Sigma}_{5}}(\mathrm{P}{\check{\Gamma}}(S))$ (resp. $\widetilde{\operatorname{Out}}^{\mathbb{I}}_{{\Sigma}_{4}}(\mathrm{P}{\check{\Gamma}}(S_{1,2}))$) in $\operatorname{Aut}(\mathrm{P}{\check{\Gamma}}(S_{0,5}))$ (resp. $\operatorname{Aut}(\mathrm{P}{\check{\Gamma}}(S_{1,2}))$).
@@
Note that $\operatorname{Aut}^{\mathbb{I}}_{{\Sigma}_{5}}(\mathrm{P}{\check{\Gamma}}(S_{0,5}))\subset\operatorname{Aut}^{\mathbb{I}}(\mathrm{P}{\check{\Gamma}}(S_{0,5}))$ and $\operatorname{Aut}^{\mathbb{I}}_{{\Sigma}_{4}}(\mathrm{P}{\check{\Gamma}}(S_{1,2}))\subset\operatorname{Aut}^{\mathbb{I}}(\mathrm{P}{\check{\Gamma}}(S_{1,2}))$. We then have:
@@
$\begin{array}[]{l}\widetilde{\operatorname{Out}}^{\mathbb{I}}(\mathrm{P}{\check{\Gamma}}(S_{1,2}))\cong\Sigma_{4}\times\widetilde{\operatorname{Out}}^{\mathbb{I}}_{{\Sigma}_{4}}(\mathrm{P}{\check{\Gamma}}(S_{1,2})),\\ \\ \widetilde{\operatorname{Out}}^{\mathbb{I}}_{{\Sigma}_{4}}(\mathrm{P}{\check{\Gamma}}(S_{1,2}))\cong\operatorname{Out}^{\mathbb{I}}(\mathrm{P}{\check{\Gamma}}(S_{1,2})),\\ \\ \operatorname{Out}^{\mathbb{I}}(\mathrm{P}{\check{\Gamma}}(S_{0,5}))\cong\Sigma_{5}\times\operatorname{Out}^{\mathbb{I}}_{{\Sigma}_{5}}(\mathrm{P}{\check{\Gamma}}(S_{0,5})).\end{array}$
@@
$\operatorname{Aut}^{\mathbb{I}}_{{\Sigma}_{4}}(\mathrm{P}{\check{\Gamma}}(S_{1,2}))\hookrightarrow\operatorname{Aut}^{\mathbb{I}}_{{\Sigma}_{5}}(\mathrm{P}{\check{\Gamma}}(S_{0,5})).$
@@
There is a short exact sequence:
@@
By [18, Corollary C] (cf. also the proof of Lemma 6.7), the image of ${\Sigma}_{5}$ in $\operatorname{Out}(\mathrm{P}{\check{\Gamma}}(S_{0,5}))$ (and so in $\operatorname{Out}^{\mathbb{I}}(\mathrm{P}{\check{\Gamma}}(S_{0,5}))$) identifies with a normal subgroup. Therefore, there is also a short exact sequence:
@@
Since $\Sigma_{4}$ and $\Sigma_{5}$ are complete groups, the above short exact sequences split (cf. [38, Theorem 7.15]) and there are natural isomorphisms as stated in the lemma. The last statement of the lemma follows from the fact that the centralizer of $\Sigma_{4}$ in $\Sigma_{5}$ is trivial. ∎
@@
By Lemma 2.19, an element $f\in\operatorname{Aut}^{\mathbb{I}}(\mathrm{P}{\check{\Gamma}}(S_{1,2}))$ preserves ${\check{\Gamma}}(S_{1,2})$-orbits. Hence, if the given $\{v_{0},v_{1}\}\in{\check{C}}_{P}(S_{1,2})_{1}$ is an edge contained in a ${\check{\Gamma}}(S_{1,2})$-orbit such that the pair of vertices $\{q(v_{0}),q(v_{1})\}$ is an edge of ${\check{C}}_{P}(S_{0,5})$, it follows that $\{\tilde{f}(q(v_{0})),\tilde{f}(q(v_{1}))\}$ is also an edge of ${\check{C}}_{P}(S_{0,5})$. In this case, we just let $w_{0}:=q(v_{0})$ and $w_{1}:=q(v_{1})$ and we are done. As it is easy to check, in particular, this happens if the common profinite simple closed curve in the intersection $v_{0}\cap v_{1}$ has the topological type of a separating curve on $S_{1,2}$.
@@
Let us then consider the case when the common profinite simple closed curve $\gamma$} in $v_{0}\cap v_{1}$ has the topological type of a nonseparating curve on $S_{1,2}$. As usual, it is not restrictive to assume that ${\gamma}\in C(S_{1,2})_{0}\subset{\check{C}}(S_{1,2})_{0}$. Moreover, by the first isomorphism of Lemma 2.21, after, possibly, composing the given $f\in\operatorname{Aut}^{\mathbb{I}}(\mathrm{P}{\check{\Gamma}}(S_{1,2}))$ with an inner automorphism of $\mathrm{P}\Gamma(S_{1,2})$, we can also assume that $f\in\operatorname{Aut}^{\mathbb{I}}_{{\Sigma}_{4}}(\mathrm{P}{\check{\Gamma}}(S_{1,2}))$.
From the last statement of Lemma 2.21 and Lemma 2.20, it then follows that the conjugacy class of ${\hat{\mathrm{I}}}_{\gamma}\cap\mathrm{P}{\check{\Gamma}}(S_{0,5})$ in $\mathrm{P}{\check{\Gamma}}(S_{0,5})$ is preserved by $\tilde{f}$. Therefore, after, possibly, composing by an inner automorphism of $\mathrm{P}{\check{\Gamma}}(S_{0,5})$, we can at last assume that $f\in\operatorname{Aut}^{\mathbb{I}}_{{\Sigma}_{4}}(\mathrm{P}{\check{\Gamma}}(S_{1,2}))_{{\hat{\mathrm{I}}}_{\gamma}}$.
@@
Since $f$ fixes the vertex $\gamma$} (cf. Remark 2.16), it also preserves the link $\operatorname{Lk}({\gamma})$ of $\gamma$} in ${\check{C}}(S_{1,2})$ and then the vertex set of the profinite subgraph $L_{\gamma}$ of ${\check{C}}_{P}(S_{1,2})$. Let us recall (cf. Lemma 2.9) that there are natural $\mathrm{P}{\check{\Gamma}}(S_{1,2})_{\gamma}$-equivariant continuous isomorphisms $\operatorname{Lk}({\gamma})\cong{\check{C}}(S_{1,2}\smallsetminus{\gamma})$ and $L_{\gamma}\cong{\check{C}}_{P}(S_{1,2}\smallsetminus{\gamma})$. By hypothesis, we then have that both $\{v_{0},v_{1}\}$ and $\{f(v_{0}),f(v_{1})\}\in(L_{\gamma})_{1}$. We claim that this implies that $f$ preserves the edge set of $L_{\gamma}$ (and so induces an automorphism of this $1$-dimensional simplicial profinite complex).
By Lemma 2.12, $f$ acts on the vertex set $(L_{\gamma})_{0}\cong{\check{C}}_{P}(S_{1,2}\smallsetminus{\gamma})_{0}$ through its image via the natural homomorphism:
@@
$R_{\gamma}\colon\thinspace\operatorname{Aut}^{\mathbb{I}}(\mathrm{P}{\check{\Gamma}}(S_{1,2}))_{{\hat{\mathrm{I}}}_{\gamma}}\to\operatorname{Aut}^{\mathbb{I}}(\mathrm{P}{\check{\Gamma}}(S_{1,2}\smallsetminus{\gamma})),$
induced by the restriction to the stabilizer $\mathrm{P}{\check{\Gamma}}(S_{1,2})_{\vec{\gamma}}$ followed by the projection to its quotient $\mathrm{P}{\check{\Gamma}}(S_{1,2}\smallsetminus{\gamma})$.
Note that $S_{1,2}\smallsetminus{\gamma}\cong S_{0,4}$. Hence, there is a short exact sequence:
(10) $1\to\mathrm{P}{\check{\Gamma}}(S_{1,2}\smallsetminus{\gamma})\to{\check{\Gamma}}(S_{1,2}\smallsetminus{\gamma})\to\Sigma_{4}\to 1$
and so a faithful representation $\Sigma_{4}\hookrightarrow\operatorname{Out}^{\mathbb{I}}(\mathrm{P}{\check{\Gamma}}(S_{1,2}\smallsetminus{\gamma}))$. Let then $\operatorname{Out}_{{\Sigma}_{4}}^{\mathbb{I}}(\mathrm{P}{\check{\Gamma}}(S_{1,2}\smallsetminus{\gamma}))$ be the centralizer of the image of ${\Sigma}_{4}$ in $\operatorname{Out}^{\mathbb{I}}(\mathrm{P}{\check{\Gamma}}(S_{1,2}\smallsetminus{\gamma}))$ and let $\operatorname{Aut}^{\mathbb{I}}_{{\Sigma}_{4}}(\mathrm{P}{\check{\Gamma}}(S_{1,2}\smallsetminus{\gamma}))$ be the inverse image of $\operatorname{Out}_{{\Sigma}_{4}}^{\mathbb{I}}(\mathrm{P}{\check{\Gamma}}(S_{1,2}\smallsetminus{\gamma}))$ in $\operatorname{Aut}^{\mathbb{I}}(\mathrm{P}{\check{\Gamma}}(S_{1,2}\smallsetminus{\gamma}))$.
Lemma 2.23.
$R_{\gamma}(\operatorname{Aut}^{\mathbb{I}}_{{\Sigma}_{4}}(\mathrm{P}{\check{\Gamma}}(S_{1,2}))_{{\hat{\mathrm{I}}}_{\gamma}})\subseteq\operatorname{Aut}^{\mathbb{I}}_{{\Sigma}_{4}}(\mathrm{P}{\check{\Gamma}}(S_{1,2}\smallsetminus{\gamma}))$.
Proof.
By Lemma 2.21, by restriction, we get a natural monomorphism:
$\operatorname{Aut}^{\mathbb{I}}_{{\Sigma}_{4}}(\mathrm{P}{\check{\Gamma}}(S_{1,2}))_{{\hat{\mathrm{I}}}_{\gamma}}\hookrightarrow\operatorname{Aut}^{\mathbb{I}}_{{\Sigma}_{5}}(\mathrm{P}{\check{\Gamma}}(S_{0,5})).$
In order to prove the lemma, we just need to translate this statement in terms of the homomorphism $R_{\gamma}$.
Let us observe that $\mathrm{P}{\check{\Gamma}}(S_{1,2})_{\vec{\gamma}}\cap\mathrm{P}{\check{\Gamma}}(S_{0,5})=\mathrm{P}{\check{\Gamma}}(S_{0,5})_{q({\gamma})}$ and that, after identifying $\mathrm{P}{\check{\Gamma}}(S_{0,5}\smallsetminus q({\gamma}))$ with $\mathrm{P}{\check{\Gamma}}(S_{0,4})$, the natural projection $\mathrm{P}{\check{\Gamma}}(S_{0,5})_{q({\gamma})}\to\mathrm{P}{\check{\Gamma}}(S_{0,5}\smallsetminus q({\gamma}))$ identifies with the restriction of a forgetful homomorphism $\mathrm{P}{\check{\Gamma}}(S_{0,5})\to\mathrm{P}{\check{\Gamma}}(S_{0,4})$ to $\mathrm{P}{\check{\Gamma}}(S_{0,5})_{q({\gamma})}$.
In [16, Section 1.2], the group $\operatorname{Aut}^{\sharp}(\mathrm{P}{\check{\Gamma}}(S_{0,n}))$, for $n\geq 4$, was defined to be the inverse image of the group $\operatorname{Out}^{\sharp}(\mathrm{P}{\check{\Gamma}}(S_{0,n}))$ by the natural homomorphism $\operatorname{Aut}(\mathrm{P}{\check{\Gamma}}(S_{0,n}))\to\operatorname{Out}(\mathrm{P}{\check{\Gamma}}(S_{0,n}))$. By Lemma 2.20, we then have that $\operatorname{Aut}^{\sharp}(\mathrm{P}{\check{\Gamma}}(S_{0,n}))=\operatorname{Aut}^{\mathbb{I}}_{{\Sigma}_{n}}(\mathrm{P}{\check{\Gamma}}(S_{0,n}))$, for $n=4,5$.
In [16, Section 2.2], it is proved that the forgetful homomorphism $\mathrm{P}{\check{\Gamma}}(S_{0,n})\to\mathrm{P}{\check{\Gamma}}(S_{0,n-1})$ induces a homomorphism $\operatorname{Aut}^{\sharp}(\mathrm{P}{\check{\Gamma}}(S_{0,n}))\to\operatorname{Aut}^{\sharp}(\mathrm{P}{\check{\Gamma}}(S_{0,n-1}))$, for $n\geq 5$.
By Lemma 2.21 and the above remarks, the restriction of the homomorphism $R_{\gamma}$ to the subgroup $\operatorname{Aut}^{\mathbb{I}}_{{\Sigma}_{4}}(\mathrm{P}{\check{\Gamma}}(S_{1,2}))_{{\hat{\mathrm{I}}}_{\gamma}}$ is then equivalent to the restriction of the homomorphism considered above $\operatorname{Aut}^{\sharp}(\mathrm{P}{\check{\Gamma}}(S_{0,5}))\to\operatorname{Aut}^{\sharp}(\mathrm{P}{\check{\Gamma}}(S_{0,4}))$ (associated to the forgetful homomorphism $\mathrm{P}{\check{\Gamma}}(S_{0,5})\to\mathrm{P}{\check{\Gamma}}(S_{0,4})$) to the image of $\operatorname{Aut}^{\mathbb{I}}_{{\Sigma}_{4}}(\mathrm{P}{\check{\Gamma}}(S_{1,2}))_{{\hat{\mathrm{I}}}_{\gamma}}$ in $\operatorname{Aut}^{\mathbb{I}}_{{\Sigma}_{5}}(\mathrm{P}{\check{\Gamma}}(S_{0,5}))=\operatorname{Aut}^{\sharp}(\mathrm{P}{\check{\Gamma}}(S_{0,5}))$. This implies the claim of the lemma. ∎
@@
By Lemma 2.23, the image of $f$ in $\operatorname{Out}^{\mathbb{I}}(\mathrm{P}{\check{\Gamma}}(S_{1,2}\smallsetminus{\gamma}))$ commutes with the image of $\Sigma_{4}$, so that all hypotheses of Lemma 2.18, applied to the short exact sequence (10), are satisfied and $f$ extends to an automorphism of ${\check{\Gamma}}(S_{1,2}\smallsetminus{\gamma})$. As in Section 2.10, we then conclude that $f$ induces an automorphism of the pants complex ${\check{C}}_{P}(S_{1,2}\smallsetminus{\gamma})$, as claimed above.
@@
It is now easy to check that, for some edge $\{v_{0}^{\prime},v_{1}^{\prime}\}\in(L_{\gamma})_{1}\cong{\check{C}}_{P}(S_{1,2}\smallsetminus{\gamma})_{1}$, we have that $\{q(v_{0}^{\prime}),q(v_{1}^{\prime})\}\in{\check{C}}_{P}(S_{0,5})_{1}$ and conclude, as we did at the beginning of the proof, letting $w_{0}:=q(v_{0}^{\prime})$ and $w_{1}:=q(v_{1}^{\prime})$. This concludes the proof of Lemma 2.22. ∎
@@
From [9, Lemma 9.13], it now follows that the normalizer of $\mathrm{P}{\check{\Gamma}}(S_{1,2})$ in ${\check{\Gamma}}^{\pm}(S_{0,5})$ coincides with the closure in this group of the normalizer of $\mathrm{P}\Gamma(S_{1,2})$ in ${\Gamma}^{\pm}(S_{0,5})$, which, as it easily follows from [28, Theorem, (ii)], is just $\mathrm{P}\Gamma^{\pm}(S_{1,2})$. Since $\operatorname{Inn}({\check{\Gamma}}^{\pm}(S_{1,2}))=\operatorname{Inn}(\mathrm{P}{\check{\Gamma}}^{\pm}(S_{1,2}))$, this implies Theorem 2.17 for $S=S_{1,2}$.